OpenStudy (anonymous):

convert 6 yards and 2 feet into inches

Vocaloid (vocaloid):

3 feet per yard, 12 inches per foot, so: 6 yards = 6*3*12 = 216 inches and 2 feet = 2*12 = 24 inches 216 + 24 = 240 inches